# Selective Excitation of Superconducting Qubits with a Shared Control Line through Pulse Shaping

### v1

@@ I Suppression of Unintended Excitation Using Selective Excitation Pulse
Figure 5 shows $P_{1}$ as a function of the pulse duration $T$. Gaussian pulses, shown as the dotted curve in Fig. 5, result in a relatively large value of $P_{1}$, indicating significant unintended excitation in the non-target qubit. In contrast, the pulse based on the SEP technique yields a much smaller $P_{1}$, demonstrating effective suppression of excitation in the non-target qubit, as shown by the solid and dot-dashed curves in Fig. 5. These results indicate that by utilizing the SEP technique, selective excitation of the target qubit can be achieved even when the frequencies of the target and non-target qubits are closely spaced.
Figure 5: Numerical calculation results for $P_{1}$, the probability of finding the non-target qubit in the excited state |$\ket{1}$ after a pulse duration of $T$. The dashed, the dot-dashed and the solid curves show the results for the Gaussian pulse, the SEP derived from Eq. (2) and the SEP obtained from Eq. (3), respectively.
II Selective Excitation Pulse for Single-Qubit Gate Operation
@@
$S(\omega)=A\left[\prod_{j\in\mathcal{Q}_{\textrm{NT}}}{(\omega-\omega_{j})(\omega-2\omega_{\textrm{d}}+\omega_{j})}\right]\exp\left[-\frac{(\omega-\omega_{\textrm{d}}-\delta)^{2}}{2\sigma^{2}}\right].$ (6)
Here, the detuning parameter $\delta$ is introduced to further fine-tune the frequency profile. By adjusting $\delta$, we compensate for residual AC-Stark shifts during gate operations [15].
The waveforms corresponding to the $X_{\pi/2}$ gate are shown in Fig. 6, and the detailed parameters are summarized in Table 2. The shaded regions in Fig. 6 represent truncated segments of the waveforms and the non-shaded regions indicate the waveforms used in the experiments.
@@
Table 2: Parameters used for Gaussian pulses and the SEPs when each qubit is the target qubit.
@@
Q0 Q1 Q2 Q0 Q1 Q2
$\delta/2\pi$ (MHz) $6.019$ $-0.3979$ $0.9885$ $6.561$ $-1.767$ $1.354$
$\Delta_{j}/2\pi$ (MHz) - - - $-77$ $-103$ $77$ $-26$ $103$ $26$
@@
III Experimental setup
The experimental setup employed in this study is illustrated in Fig. 7. The device consisting of 64 fixed-frequency transmon qubits was mounted to the stage of a dilution refrigerator (Bluefors XLD1000sl) at approximately 10 mK and was magnetically shielded.
@@ III Experimental setup
The qubit readout was performed using a dispersive readout technique [28]. The readout resonator in the device was dispersively coupled to a set of four transmon qubits, allowing for the simultaneous readout of all four qubits. In this setup, Q1 and Q2 shared the same measurement line. The readout signals were amplified by high-electron-mobility transistor (HEMT) amplifiers located at the 4K stage before being routed back to the qubit controllers.
Figure 7: Wiring schematic of experimental setup. Background color indicates the temperature at each stage. The setup allows pulses driving the target qubit to interact with non-target qubits via a signal divider.

### v2

@@ 1-2 Machikaneyama, Toyonaka, Osaka, 560-0043, Japan
[…]
$S_{\mathrm{asym}}(\omega)=A(\omega-\omega_{1})\exp\left[-\frac{(\omega-\omega_{\textrm{d}})^{2}}{2\sigma^{2}}\right],$ (S2)
and
$S_{\mathrm{sym}}(\omega)=A(\omega-\omega_{1})(\omega-2\omega_{\textrm{d}}+\omega_{1})\exp\left[-\frac{(\omega-\omega_{\textrm{d}})^{2}}{2\sigma^{2}}\right].$ (S3)
Here, $A$ is the amplitude and $\sigma$ determines the width of the Gaussian envelope. The parameters used for this calculation are $\sigma/2\pi=1/T$. The amplitude $A$ and the drive frequency $\omega_{\textrm{d}}$ are chosen to satisfy $\int_{0}^{T}s_{\mathrm{asym\,(sym)}}(t)\,dt=\pi/2$ and $\Delta_{0}=0$, where $s_{\mathrm{asym\,(sym)}}(t)$ is respectively obtained by applying the inverse Fourier transform to $S_{\mathrm{asym\,(sym)}}(\omega)$, followed by truncation within the pulse duration $T$.
Figure S1 summarizes $P_{\textrm{e}}$ as functions of the pulse duration $T$ and the detuning $\Delta_{1}$. Figures S1(a)–(c) show the excitation probability of the target qubit Q0 driven by different pulses: (a) Gaussian pulse, (b) SEP $s_{\mathrm{asym}}(t)$, and (c) SEP $s_{\mathrm{sym}}(t)$. Figures S1(d)–(f) show the corresponding excitation probability $P_{\textrm{e}}$ for the non-target qubit Q1. For the target qubit Q0, the Gaussian pulse [Fig. S1(a)] achieves $P_{\textrm{e}}>0.49$ with relatively short durations (approximately 20 ns), while SEP-based pulses [Figs. S1(b) and (c)] generally require longer durations to reach similar excitation levels. In contrast, for the non-target qubit Q1, SEP-based pulses [Figs. S1(e) and (f)] provide significantly better suppression of unwanted excitation than the Gaussian pulse [Fig. S1(d)], particularly when the detuning $\Delta_{1}$ is small. These results indicate that by utilizing the SEP technique, selective excitation of the target qubit can be achieved even when the frequencies of the target and non-target qubits are closely spaced.
Figure S1: Numerical simulation of the excitation probability $P_{\textrm{e}}$ of the target qubit Q0 and the non-target qubit Q1 after an $X_{\pi/2}$ gate, shown as a function of the pulse duration $T$ and detuning $\Delta_{1}$. (a)–(c) Results for Q0 using three waveform types: (a) Gaussian, (b) $s_{\mathrm{asym}}(t)$, and (c) $s_{\mathrm{sym}}(t)$. (d)–(f) Results for Q1, using the same waveform types as in (a)–(c).
S2 Selective-Excitation Pulse for Single-Qubit Gate Operation
@@
$S(\omega)=A\left[\prod_{j\in\mathcal{Q}_{\textrm{NT}}}{(\omega-\omega_{j})(\omega-2\omega_{\textrm{d}}+\omega_{j})}\right]\exp\left[-\frac{(\omega-\omega_{\textrm{d}}-\delta)^{2}}{2\sigma^{2}}\right].$ (S4)
Here, the detuning parameter $\delta$ is introduced to further fine-tune the frequency profile. By adjusting $\delta$, we compensate for residual ac Stark shifts during gate operations [15].
The waveforms corresponding to the $X_{\pi/2}$ gate are shown in Fig. S2, and the detailed parameters are summarized in Table S1.
@@
Table S1: Parameters used for the Gaussian pulses and the SEPs.
@@
Target qubit Q0 Q1 Q2 Q0 Q1 Q2
$\delta/2\pi$ (MHz) $6.019$ 0.3979-0.3979 $0.9885$ $6.561$ 1.767-1.767 $1.354$
$\Delta_{j}/2\pi$ (MHz) 77-77 103-103 $77$ 26-26 $103$ $26$
@@
S3 Discontinuities in the Selective Excitation Pulse
Due to the finite temporal duration of the waveform, SEPs exhibit discontinuities at their start and end points, as shown in Fig. S2. These discontinuities are expected to introduce significant spectral leakage, particularly for shorter pulses where truncation effects become more pronounced. To mitigate this issue and improve waveform smoothness, we investigate the role of the spectral width parameter $\sigma$ in Eq. (S3), which defines the envelope width of the frequency-domain profile. The amplitude $A$ is normalized such that the time-domain waveform $s(t)$, obtained via inverse Fourier transform of $S(\omega)$, satisfies the area condition $\int^{T}_{0}s(t)\,dt=\pi/2$, with a fixed truncated pulse duration $T=40$ ns. We set $\Delta_{1}/2\pi=0.03$ GHz, where $\Delta_{1}$ denotes the detuning between the drive frequency $\omega_{\mathrm{d}}$ and the non-target frequency $\omega_{1}$ in Eq. (S3).
[…]
S4 Experimental setup
The experimental setup employed in this study is illustrated in Fig. S4. The device consisting of 64 fixed-frequency transmon qubits was mounted to the stage of a dilution refrigerator (Bluefors XLD1000sl) at approximately 10 mK and was magnetically shielded.
@@
The qubit readout was performed using a dispersive readout technique [29]. The readout resonator in the device was dispersively coupled to one of the four transmon qubits in a unit cell, allowing for the simultaneous readout of all four qubits. In this setup, Q1 and Q2 shared the same measurement line. The readout signals were amplified by high-electron-mobility transistor (HEMT) amplifiers located at the 4-K stage before being routed back to the qubit controllers.
Figure S4: Wiring schematic of the experimental setup. Background color indicates the temperature at each stage. The setup allows pulses driving the target qubit to interact with non-target qubits via a signal divider.